OpenStudy (anonymous):

Which of the following points lies on the graph of the function y = 8x? The Choices are... 1. (0,1) 2. (8,1) 3. (2,64) 4. (0,8) Please Help Me! <:v

OpenStudy (anonymous):

have you tried substituting the points into the function?

OpenStudy (anonymous):

oh i meant it 8^x

OpenStudy (anonymous):

put the values of and get y and compare with ans

OpenStudy (anonymous):

srry :/

OpenStudy (anonymous):

remember that coordinates are always in the form (x,y) so if we look at the first choice x=0 and y=1 we substitute those points into the function y=8^x 1=8^0 if it doesnt make sense, then the point does not lie on the function if it does make sense, then that point lies on your function

OpenStudy (anonymous):

now check the rest of the points

OpenStudy (anonymous):

ohokie thanks :D